# MCTS-RAG: Enhance Retrieval-Augmented Generation with Monte Carlo Tree Search

### v1

@@ 4.3 Fine-grained Analysis
We evaluate the effectiveness of retrieval actions and rollout times in subsection 4.1. Specifically, we conduct an ablation by disabling different retrieval modules (A4, A5, or both) to gauge their impact on overall performance. In addition, we vary the number of rollouts from 4 to 16 to investigate how deeper search affects accuracy and efficiency.
@@
Retrieval actions, especially A4 and A5, are key for multi-step reasoning. Enabling all retrievals boosts GPQA (+20.88%) and FMT (+16.10%). Disabling A5 improves GPQA (+7.84%) and FMT (+6.43%) over disabling A4, suggesting A4’s stronger role. CWQA sees minimal impact (+1.25%). These findings highlight retrieval trade-offs and the importance of recursive evidence aggregation.
@@
More rollouts enhance performance, particularly for GPQA. Increasing from 4 to 8 slightly aids CWQA (+3%), while 8 to 12 boosts GPQA (+11%). Scaling to 16 further improves GPQA (+9%) and FMT (+5%), reinforcing the value of iterative reasoning.
@@ 4.4 Human Analysis and Case Study
Our case study reveals the following two key improvements introduced by MCTS-RAG: (1) Enhanced External Knowledge Utilization: Compared to other reasoning methods, MCTS-RAG achieves higher accuracy, primarily due to its richer reasoning space and more effective utilization of external knowledge. Figure 6 clearly illustrates how Monte Carlo Tree Search tightly integrates reasoning and retrieval processes, significantly enhancing the quality and richness of information used during reasoning, thereby substantially improving inference accuracy. (2) Reduced Hallucination Risks: Moreover, MCTS-RAG mitigates hallucination risks through detailed and explicit reasoning steps. On one hand, the explicit reasoning pathways enable the model to more accurately interpret retrieved external knowledge, reducing errors arising from ambiguity or misunderstanding (as illustrated in Figure 7 in Appendix). On the other hand, these thorough reasoning procedures generate clearer and more contextually relevant queries, thus improving the precision of external information retrieval (as illustrated in Figure 9 in Appendix). Consequently, MCTS-RAG demonstrates substantial advantages over traditional reasoning methods in terms of improved accuracy and robustness.
Error Case Analysis
@@ 5 Conclusion
In this work, we propose MCTS-RAG, an approach that integrates Monte Carlo Tree Search with Retrieval-Augmented Generation to improve multi-step reasoning accuracy and reliability. By effectively allocating search and retrieval processes, MCTS-RAG excels at handling cross-domain tasks that require in-depth external knowledge. It not only enables flexible formulation of high-quality retrieval queries but also refines the reasoning path through iterative tree exploration, thus reducing hallucinations caused by shallow retrieval or simplistic reasoning. Experimental results show that MCTS-RAG has achieved good results in scenarios such as complex reasoning tasks, knowledge-enhanced scientific question-answering tasks, and challenging fact-checking tasks.
Acknowledgments
We are grateful to Google TRC program for providing computing resources and Together AI for granting LLM API credits.
@@ Acknowledgments
MCTS-RAG integrates MCTS-based reasoning and RAG to enhance reasoning capabilities, but several errors persist. Amplification errors occur when early retrieval mistakes propagate through search iterations. Factual confusion arises from semantic mismatches leading to incorrect reasoning. Information overload happens when excessive retrieval results cause reasoning to deviate from the target. Additionally, search latency remains a challenge, as deep MCTS search trees significantly increase reasoning time, particularly with multiple retrieval steps. Action selection complexity arises because the optimal choice among A1-A6 depends on query difficulty, necessitating a more adaptive decision mechanism. Inefficient expansion occurs when MCTS explores unnecessary branches due to a lack of effective pruning based on retrieval confidence or early error detection. Addressing these issues is essential for improving efficiency and reasoning accuracy.
We encourage future work to focus on optimizing search efficiency by developing adaptive action selection strategies, confidence-based retrieval filtering, and error-aware pruning mechanisms to improve MCTS exploration. Additionally, integrating reinforcement learning for dynamic search policy refinement may further enhance reasoning accuracy. Addressing these challenges will contribute to the development of more robust and scalable reasoning models, bridging the gap between retrieval-based methods and human-like problem-solving.
@@ Appendix A Prompts for Each Action
A chat between a curious user and an AI assistant. The assistant gives step-by-step solutions to the user’s questions. In the end of the assistant’s response, a final answer must be given in the format of "The answer is: <ANSWER>.", where <ANSWER> should be a concise answer. Usage Example: {examples} Instruction: {instruction} Note: Please answer in a complete sentence.
@@
Template: A chat between a curious user and an AI assistant. The assistant gives step-by-step solutions to the user’s questions with each step numbered. At the final step, a conclusive answer must be given in the format "The answer is: <ANSWER>.", where <ANSWER> should be a concise answer. Instruction: {instruction} Note: Let’s think step by step.
@@
Template: Given a question, decompose it into sub-questions. For each sub-question, provide an answer in one complete sentence ending with "The answer is ". When the original question is answerable, start the sub-question with "Now we can answer the question: <original question>".
@@
Template: Given a question, generate a search query that would help gather information to answer it. Your goal is to formulate a query that retrieves useful evidence or additional details relevant to the question. The query should be specific enough to ensure that the search results are both relevant and helpful. Please answer in one complete sentence, starting with "The query is: <your retrieve query>". Question: {question}
@@
Template: A chat between a curious user and an AI assistant. The assistant evaluates whether the retrieved information is relevant to the search query and sufficient to answer the question. Please provide a concise evaluation in one complete sentence, starting with "Evaluation:". Instruction: Please assess if the retrieved information is related to the query and can be used to answer the question.
@@
Template: Analyze the provided Knowledge and extract key information relevant to the Original Question. Present your analysis in a concise and organized format. Input: - Original Question: {original_question} - Knowledge: {retrieved_context} Output Format: Key Points: Point 1: Relevant information; Point 2: Relevant information; Point 3: Relevant information... Requirement: The output must be a single line summarizing all key points in one sentence without redundant description.

### v2

@@ 4.6 Fine-grained Analysis
We conduct an ablation by disabling different retrieval modules (A4, A5, or both) to gauge their impact on overall performance. In addition, we vary the number of rollouts from 4 to 16 to investigate how deeper search affects accuracy and efficiency. Table 2 shows the results.
@@
Retrieval actions, especially A4 and A5, are key for multi-step reasoning. Enabling all retrievals boosts GPQA (+32.3%) and FMT (+17.9%). Disabling A5 improves GPQA (+11.8%) and FMT (+12.0%) over disabling A4, suggesting A4’s stronger role. CWQA sees minimal impact (+5.9%). These findings highlight retrieval trade-offs and the importance of recursive evidence aggregation.
@@
More rollouts enhance performance. Specifically, increasing from 4 to 8 slightly aids CWQA (+3.0%), while 8 to 12 boosts GPQA (+11.5%). Scaling to 16 further improves GPQA (+9.1%) and FMT (+4.7%), showing the value of iterative reasoning.
@@ 4.7 Human Analysis and Case Study
Our case study reveals the following two key improvements introduced by MCTS-RAG: (1) Enhanced External Knowledge Utilization: Compared to other reasoning methods, MCTS-RAG achieves higher accuracy, primarily due to its richer reasoning space and more effective utilization of external knowledge. Figure 8 clearly illustrates how Monte Carlo Tree Search tightly integrates reasoning and retrieval processes, significantly enhancing the quality and richness of information used during reasoning, thereby substantially improving inference accuracy. (2) Reduced Hallucination Risks: Moreover, MCTS-RAG mitigates hallucination risks through detailed and explicit reasoning steps. On one hand, the explicit reasoning pathways enable the model to more accurately interpret retrieved external knowledge, reducing errors arising from ambiguity or misunderstanding (as illustrated in Figure 9 in Appendix). On the other hand, these thorough reasoning procedures generate clearer and more contextually relevant queries, thus improving the precision of external information retrieval (as illustrated in Figure 11 in Appendix).
Error Case Analysis.
@@ 5 Conclusion
The work introduces MCTS-RAG, which integrates MCTS with RAG to improve multi-step reasoning accuracy and reliability. MCTS-RAG not only enables flexible formulation of high-quality retrieval queries but also refines the reasoning path through iterative tree exploration, thus reducing hallucinations caused by shallow retrieval or simplistic reasoning. To further enhance practicality, we introduce several efficiency optimizations, including parallel expansion and retrieval pruning, significantly reducing inference latency without compromising accuracy. Experimental results demonstrate that MCTS-RAG achieves strong performance on complex reasoning, knowledge-enhanced scientific QA, and fact-checking tasks.
@@
MCTS-RAG integrates MCTS-based reasoning and RAG to enhance reasoning capabilities, but several errors persist. Amplification errors occur when early retrieval mistakes propagate through search iterations. Factual confusion arises from semantic mismatches leading to incorrect reasoning. Information overload happens when excessive retrieval results cause reasoning to deviate from the target. Additionally, search latency remains a challenge, as deep MCTS search trees significantly increase reasoning time, particularly with multiple retrieval steps. Action selection complexity arises because the optimal choice among A1-A6 depends on query difficulty, necessitating a more adaptive decision mechanism. While our current heuristic-based action space design provides a practical balance between task complexity and potential reward, it may not optimally adapt to all query types. We further acknowledge that dynamically adjusting the action space could improve flexibility in more complex environments, but may also introduce hallucinations or instability, particularly for small-capacity models. Inefficient expansion occurs when MCTS explores unnecessary branches due to a lack of effective pruning based on retrieval confidence or early error detection. Addressing these issues is essential for improving efficiency and reasoning accuracy.
We encourage future work to focus on optimizing search efficiency by developing adaptive action selection strategies, confidence-based retrieval filtering, and error-aware pruning mechanisms to improve MCTS exploration. Additionally, integrating reinforcement learning for dynamic search policy refinement may further enhance reasoning accuracy. Exploring ways to refine dynamic action-space adjustment while mitigating instability for small models is another promising direction. Addressing these challenges will contribute to the development of more robust and scalable reasoning models, bridging the gap between retrieval-based methods and human-like problem-solving.
@@ Appendix A Prompts for Each Action
Template:
A chat between a curious user and an AI assistant. The assistant gives step-by-step solutions to the user’s questions. In the end of the assistant’s response, a final answer must be given in the format of "The answer is: <ANSWER>.", where <ANSWER> should be a concise answer.
Usage Example:
{examples}
Instruction:
{instruction}
Note:
Please answer in a complete sentence.
@@
Template:
A chat between a curious user and an AI assistant. The assistant gives step-by-step solutions to the user’s questions with each step numbered. At the final step, a conclusive answer must be given in the format "The answer is: <ANSWER>.", where <ANSWER> should be a concise answer.
Instruction:
{instruction}
Note:
Let’s think step by step.
@@
Template:
Given a question, decompose it into sub-questions. For each sub-question, provide an answer in one complete sentence ending with "The answer is ". When the original question is answerable, start the sub-question with "Now we can answer the question: <original question>".
@@
Template:
Given a question, generate a search query that would help gather information to answer it. Your goal is to formulate a query that retrieves useful evidence or additional details relevant to the question. The query should be specific enough to ensure that the search results are both relevant and helpful. Please answer in one complete sentence, starting with "The query is: <your retrieve query>".
Question:
{question}
@@
Template:
A chat between a curious user and an AI assistant. The assistant evaluates whether the retrieved information is relevant to the search query and sufficient to answer the question. Please provide a concise evaluation in one complete sentence, starting with "Evaluation:".
Instruction:
Please assess if the retrieved information is related to the query and can be used to answer the question.
@@
Template:
Analyze the provided Knowledge and extract key information relevant to the Original Question. Present your analysis in a concise and organized format.
Input:
- Original Question: {original_question}
- Knowledge: {retrieved_context}
Output Format:
Key Points: Point 1: Relevant information; Point 2: Relevant information; Point 3: Relevant information...
Requirement:
The output must be a single line summarizing all key points in one sentence without redundant description.